OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] Plug the point in this formula to get the equation. As for perpendicular slopes, the rule is: \[m _{1}=-\frac{ 1 }{ m _{2} }\]

OpenStudy (anonymous):

@Ankh 2x+y= 13? I got partial credit though..

OpenStudy (anonymous):

The perpendicular slope to -2 is 1/2 y-3=0.5(x-5) y-3=0.5x-2.5 y=0.5x-2.5+3 y=0.5x+0.5